# Machine Unlearning: A Comprehensive Survey

### v1

@@ 10. Summary
The survey aims to offer a comprehensive and up-to-date overview of machine unlearning techniques in a systematic way. We organize the main challenges, research advancements, corresponding techniques, and privacy and security issues in machine unlearning. Additionally, we presented a detailed and unified classification of machine unlearning. We first illustrate the complete unlearning framework, including the learning, request, unlearning, and verification. Then, we catalogue recent studies briefly into exact unlearning and approximate unlearning, and we introduce the technical details of these unlearning methods. Moreover, we noticed some new unlearning scenarios, such as federated unlearning and graph unlearning, which are also introduced in the survey. Besides, we consider privacy and security issues in machine unlearning as an important consisting part. We collect and summarize the related literature about unlearning privacy threats and applications. Ultimately, the survey provides clear summaries and comparisons between various unlearning scenarios and corresponding methods, giving a comprehensive view of current work and listing the challenges and open problems of different scenarios of machine unlearning.
We hope our survey can help classify future unlearning studies, achieve a more in-depth understanding of unlearning methods, and address complex challenges. We believe the open problems listed in Section 9 will still be challenging in the following years, and we will try to optimize some of them. Last but not least, we expect this survey can help researchers in the study of machine unlearning, regardless of the unlearning strategies or unlearning privacy and security threats or applications of unlearning.

### v2

@@ 10. Summary
The survey aims to offer a comprehensive and systematic overview of machine unlearning techniques. We organize the main challenges, research advancements, corresponding techniques, and privacy and security issues in machine unlearning. Additionally, we presented a detailed and unified classification of machine unlearning. We first illustrate the complete unlearning framework, including the learning, request, unlearning, and verification. Then, we briefly categorize recent studies into exact and approximate unlearning and introduce the technical details of corresponding unlearning methods. Moreover, we noticed some new unlearning scenarios, such as federated unlearning and graph unlearning, which are also introduced in the survey. Besides, we consider privacy and security issues in machine unlearning to be an important part of the studies. We collect and summarize the related literature about unlearning privacy threats and applications. Ultimately, the survey provides clear summaries and comparisons between various unlearning scenarios and corresponding methods, giving a comprehensive picture of existing work and listing the challenges and open problems of different scenarios of machine unlearning.
We hope our survey can help classify future unlearning studies, achieve a more in-depth understanding of unlearning methods, and address complex challenges. We believe the open problems listed in Section 9 will still be challenging in the following years, and we will try to optimize some of them. Last but not least, we expect the survey can help researchers in the study of machine unlearning, regardless of the unlearning strategies or unlearning privacy and security threats or applications of unlearning.